Camp Rainbow offers overnight summer camp programs for children ages 10-14 every summer during June and July. Each camp session is one week and can accommodate up to 200 children. The camp is not coed, so boys attend during the odd-numbered weeks and girls attend during the even-numbered weeks. While at the camp, participants make crafts, participate in various sports, help care for the camps resident animals, have cookouts and hayrides, and help assemble toys for local underprivileged children The camp provides all food as well as materials for all craft classes and the toys to be assembled. One cabin can accommodate up to 10 children, and one camp counselor is assigned to each cabin. Three camp managers are on-site regardless of the number of campers enrolled. Following is the cost information for Camp Rainbows operations last summer: Number of Cost to Run Week Campers 94 140 108 120 118 100 90 124 $8,630 11,740 8,830 9,400 9,010 8,680 9,590 Required 1. Perform a least-squares regression analysis on Camp Rainbows data. (Use Microsoft Excel or a statistical package to find the coefficients using least-squares regression. Round your answers to 2 decimal places.) Coefficients Intercept X Variable 1

2. Using the regression output, create a cost equation (y a+bx) for estimating Camp Rainbows operating costs. Total Cost- Number of Campers 3. Using the least-squares regression results, calculate the camps expected operating cost if 110 children attend a session. (Round your intermediate calculations and final answer to 2 decimal places.) tal Cost References eBook &Resources Use least-squares regression to estimate cost behavior.
Answer #1

Following is the output of regression analysis:

SUMMARY OUTPUT Regression Statistics Multiple R R Square Adjusted R Square Standard Error Observations 0.863309048 0.745302513 0.702852932 570.2718262 8 ANOVA df MS Significance F Regression Residual Total 1 5709827.7665709827.76617.55735845 325209.9557 0.005748295 1951259.734 7661087.5 7 Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Intercept Number of campers, X 3333.563017 53.49160611 1440.781584 12.76603557 2.313718509 0.059962599-191.9025082 0.005748295 6859.028542 84.72896978 4.190150171 22.25424245

1;

Intercept: 3333.56

X variable 1: 53.49

2:

The regression line is

y' = 3333.56 + 53.49x

3:

The expected cost is

y' = 3333.56 + 53.49*110 = 9217.46